OpenStudy (anonymous):

Help!!!! I'll Medal!!!!!! Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximatly how long does it take until the soccer ball hits the ground again?

OpenStudy (anonymous):

You Think?

OpenStudy (anonymous):

12 seconds

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

its probably 14

OpenStudy (anonymous):

6seconds???????????

OpenStudy (anonymous):

Most Likely 6 seconds But It Could be shorter are there any answer choices?

OpenStudy (anonymous):

let h be the maximum height. initial velocity u=25 ft/sec final velocity v=0 ft/sec g=-32 ft/sec \[v=u+g t\] 0=25-35t 35 t=25 \[t=\frac{ 25 }{ 32 } \sec\] Total time taken to reach the ground again \[=2\times \frac{ 25 }{ 32 }=25/16 \sec.\]